It has been found that 26% of men 20 years and older suffer from hypertension (high blood pressure) and 31.5% of women are hypertensive. A random sample 150 of each gender was selected from recent hospital records, and the following results were obtained.

Men. 43 patients had high blood pressure

Woman. 52 patients had high blood pressure.

(a) Construct 95% confidence interval for the difference of the two proportion. Round your answer to nearest ten-thousandth. Interpret the result.

(b) Based on 95% confidence interval, does a higher percentage of women have high blood pressure?  Explain.
Answer #1

Given

\small p_1 = \frac{43}{150} =0.28667

\small p_2 = \frac{52}{150} =0.34667

a)

we know that the confidence interval for the difference in proportion is given by,

\small Confidence \;interval = (p_1 - p_2)\pm z * \sqrt{\frac{p_1(1-p_1)}{n_1} +\frac{p_2(1-p_2)}{n_2} }\small Confidence \;interval = (0.28667 - 0.34667)\pm 1.96 * \sqrt{\frac{0.28667(1-0.28667)}{150} +\frac{0.34667(1-0.34667)}{150} }

\small Confidence \;interval = [-0.1482\; , \; 0.0282]

b)

Since the 95% confidence interval includes the value 0. the difference between the two proportions might be zero. Therefore, we cannot conclude that a higher percentage of women have high blood pressure.

No. Based on 95% confidence interval, a higher percentage of women do not have high blood pressure.